OpenStudy (anonymous):

a ball is thrown vertically upward from the ground its distance in feet from the ground in t seconds is s=-16t^2+144t. After how many seconds will the ball be 288 feet from the group?

OpenStudy (jack1):

\(s =-16t^2+144t\) and s =288 \(288 =-16t^2+144t\) so turn this into a quadratic equation, then solve for t \( 0 =-16t^2+144t - 288\) now solve this using the quadratic formula ... or factorize \(0 = 16( t-3 ) (t-6 )\) so t = 3 and t = 6 is where the ball will be 288 meters high